Helium is expensive so when it boils off the superconducting magnet in magnetic resonance
imaging (MRI) scanners, it is collected and recycled. As a part of the recycling process, the
monatomic He gas is first isothermally compressed by a factor of 4, then isochorically cooled to
140 K. Suppose we start with 40.0 L of the gas in the collecting bag at room temperature (300 K)
and atmospheric pressure.
The number of moles of the gas is closest to :
A. 0.8
B. 1.6
C. 2.4
D. 3.2
E. 4.0
1
Expert's answer
2018-06-19T08:52:08-0400
n(He) = V/VM = 40L/22.4L/mol = 1.79 mol
The number of moles of the gas is closest to :
B. 1.6